Convert each part of the ratio to percentage: 1: 2: 5

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have to find total parts by adding the given ratio = 1 + 2 + 5 = 8
1stpart = 1/8 = (1/8) × 100 %
= (100/8) %
= 12.5%
2ndpart = 2/8
= (2/8) × 100%
= (200/8)
= 25%
3rdpart = 5/8
= (5/8) × 100%
= (500/8) = 62.5%

Thanks